OpenStudy (anonymous):

Use the rule of differentiation to find the derivative of the function? h(x)=10/7(x)^2

OpenStudy (bahrom7893):

Did you mean: \[h(x) = 10/(7x^2)\] ?

OpenStudy (anonymous):

Squareroot just in (x)^2 10/7(x)^2

OpenStudy (bahrom7893):

Square root? Or square? And I didn't mean that I meant is x^2 in the denominator too?

OpenStudy (anonymous):

-10/7x

OpenStudy (bahrom7893):

If it's 10/7(x)^2, then it is h(x) = (10/7)x^(-2) h'(x) = (-20/7)x^(-3)

OpenStudy (anonymous):

Thank you so much.

OpenStudy (bahrom7893):

Fan us if we helped, thanks.

OpenStudy (anonymous):

\[5\sin \alpha/3-2\alpha\]

OpenStudy (bahrom7893):

(5/3)Cosa - 2

OpenStudy (anonymous):

5/3) Cos(a/3) - 2

OpenStudy (bahrom7893):

wait is it Sin(a/3) or (Sina)/3, ppl use parenthesis!!